????? ling Learning The table below presents information on the hypothetical closed economy. Use the table to answer the 1. Enter he value of private savings in 2012 2010 2011 S112 S133 $82 $22 2012 $145 $86 S26 S33 S39 Number Private savings is income (GDP) left over after GDP$ billion taxes and consumption in billions) in billions) S 20 spending with transfer paymenis added back in. S75 2. Enter the value of public savings in 2012 n billions) in billions) in bilions) S15 S22 TaxesS28 Number S-13 billion Transter 3. Enter the value of national savings in 2012 S15, $19 n bilions) billien
Answer #1

1. Private saving = GDP - C - T + TR = 145 - 86 - 39 + 19 = 39 billion

3. National saving = Private saving + Public saving = 39 billion - 13 billion = 26 billion